OpenStudy (anonymous):

a horse race has 11 entries assuming that there are no ties what is the proability that 5 horses own by one person finish first , second third and forth please help me solve

OpenStudy (anonymous):

help anyone

OpenStudy (kropot72):

The probability is zero, the reason being 5 horses cannot occupy 4 places without ties

OpenStudy (across):

I think he meant "and so forth."

OpenStudy (kropot72):

@across you could be right, but I assumed "...fourth"

OpenStudy (dumbcow):

\[P = \frac{5}{11}*\frac{4}{10}*\frac{3}{9}*\frac{2}{8}*\frac{1}{7}\]